OpenStudy (anonymous):

A cup of coffee is served at 150 degrees Fahrenheit and is left in a room that is 70 degrees Fahrenheit. Suppose that when the coffee is first placed in the room it is cooling at a rate of 20 degrees per minute. How long does it take for the coffee to cool to 100 degrees? A. 2 ½ minutes B. 3 minutes C. 4 minutes D. 4 ½ minutes

OpenStudy (perl):

you can use newtons law of cooling here

OpenStudy (anonymous):

A

OpenStudy (anonymous):

fan and medal because 20*2.5 is 50 150-50=100 @kiker

OpenStudy (anonymous):

so wait how is it a?

OpenStudy (perl):

this is really a differential equation problem. Let T(t) = temperature of object at time t T(0)= T_o =initial temperature of object T_a = ambient temperature

OpenStudy (perl):

\[\frac{dT}{dt} = -k(T-T_a)\]

OpenStudy (perl):

This has the solution \[T(t)=T_a + (T_o-T_a)e^{-kt} \]

OpenStudy (sleepyhead314):

I'm learning this too would it be ln(3/8) / 20 = 0.049 = 2.9 minutes so would it be B?

OpenStudy (perl):

\[T(t)=T_a + (T_o-T_a)e^{-kt} \\ \text{this gives us} \\ T(t) = 70 + 80e^{-kt}\]

OpenStudy (perl):

ok so far?

OpenStudy (sleepyhead314):

they gave cooling at 20 degrees per minute...

OpenStudy (sleepyhead314):

ohhh

OpenStudy (sleepyhead314):

ln(3/4) = 0.28768 = k?

OpenStudy (perl):

there is a problem very similar to this one here: https://math.dartmouth.edu/~klbooksite/3.02/302examples/302coffee.htm

OpenStudy (perl):

just change the 170 degrees to 150 degrees

OpenStudy (perl):

I would do the problem over again, but i am experiencing terrible lag.